# Boltzmann Bridges

### v1

@@ Abstract
It is often stated that the second law of thermodynamics follows from the condition that at some given time in the past the entropy was lower than it is now. Formally, this condition is the statement that $E[S(t)|S(t_{0})]$, the expected entropy of the universe at the current time $t$ conditioned on its value $S(t_{0})$ at a time $t_{0}$ in the past, is an increasing function of $t$. We point out that in general this is incorrect. The epistemic axioms underlying probability theory say that we should condition expectations on all that we know, and on nothing that we do not know. Arguably, we know the value of the universe’s entropy at the present time $t$ at least as well as its value at a time in the past, $t_{0}$. However, as we show here, conditioning expected entropy on its value at two times rather than one radically changes its dynamics, resulting in a unexpected, very rich structure. For example, the expectation value conditioned on two times can have a maximum at an intermediate time between $t_{0}$ and $t$, i.e., in our past. Moreover, it can have a negative rather than positive time derivative at the present. In such “Boltzmann bridge” situations, the second law would not hold at the present time. We illustrate and investigate these phenomena for a random walk model and an idealized gas model, and briefly discuss the role of Boltzmann bridges in our universe.
@@ I Introduction
The formal justification of the second law of thermodynamics despite the time-symmetry of the laws of physics has been a major subject of debate for well over a century. A common claim is that the second law can be understood as a simple consequence of the fact that past entropy was lower than it is now, despite the time-symmetry of Boltzmann’s H-theorem. Here we point out that this statement can be wrong, and so is misleading without careful qualifications. Sparsely in the literature is the second law or H-theorem properly written down formally in the language of stochastic variables; we do so here. Specifically, consider a statistical system, with an entropy $S(t)$ evolving in time. Boltzmann’s H-theorem can be summarized as the statement that the conditional expected value of future entropy given current entropy is an increasing function, i.e., for all $k>0$,
@@
$\frac{d}{dt}\mathbb{E}\left[S(t)\,|\,S(t_{0})\right]>0$ (3)
@@
Accordingly, if we are given the fact that in our universe, $S(t_{0})$ for a time in our past was sufficiently smaller than its maximal value, then the time-derivative of the expected entropy of the universe for all later times, conditioned on that previous value of entropy and nothing else, is non-negative. This is a fully formal statement of a common justification that has been offered for the second law, that is follows from a “past hypothesis” that entropy was particularly low some time in our past [1, 2, 3, 4].
@@
In this paper we investigate the $t$-dependence of the two-time conditioned expected value of the entropy. We begin by showing how to express the two-time conditioned expected entropy in terms of two different one-time conditioned expected entropies. This gives a particularly simple formula for two-time conditioned expected entropy whenever the one-time conditioned expected entropy is time-symmetric — as it is in the time-symmetric version of Boltzmann’s H-theorem in Eq. (2).
We then apply this result to analyze the two-time conditioned expected entropy in toy models of the dynamics of the entropy of the universe. The first of these is a random walk model for the evolution of the state of the universe, and the second is a gas randomly diffusing between two boxes, with both entropies based on Boltzmann’s.
We define “2nd law violating” curves (2LV) as solutions to $S(t)$ which have a maximum before $t_{f}$. Intuitively, the 2LV phenomenon can be seen to be true for instance if $t$ is longer than the relaxation time of the system, or if $S_{now}-S_{past}$ is less that the expected value of the entropy increase (given $S_{past}$ alone) in the $t$ interval.
These observations suggest that the interpretation of the second law as a simple consequence of a “past hypothesis” (an assumption about past entropy) requires, at the very minimum, additional careful qualifications. If we use probability theory properly, conditioning on what we know and nothing else, then the past hypothesis that $S(t_{0})$ was low does not establish that the second law holds in the present. Indeed, when probability theory is used correctly, a wealth of rich mathematics is revealed. In particular, we show that both models have the potential to exhibit changes in the concavity and derivative of $\mathbb{E}\left[S(t)\,|\,S(t_{0}),S(t_{f})\right]$, in stark contrast to the familiar behavior. In the first model, this change in concavity common among the space of model parameters. However, in the second model, there is a stronger regularity between 2LV and non-2LV curves that might allow for the computation of approximate conditions on the initial and final conditions and $t_{f}-t_{0}$ that delineate between 2LV and non-2LV behavior. Both models also exhibit unexpected behavior where the two-time conditioned entropy can cross the relaxation curve.
In Sec. VI, we discuss the relevance of these observation for our understanding of the nature of the second law in our universe. We also outline open questions about the framework and models in addition to raising questions about cosmology.
@@ II From physics to distributions
We start from a discrete-valued, discrete time Markov process $x$ taking values on a real set $I$ with a time-symmetric and time-translation invariant one-time conditioned distribution:
@@
The general set-up consists of computing the conditional probability of the stochastic process $x$ conditioned on 2 times: $\mathbf{P}(x(t)\,|\,x(t_{0}),x(t_{f})),$ with $t_{0}<t<t_{f}$. It is critical to the problem that we seek time-dependence of expectation values in between the initial and final conditions, otherwise one could simply use Markovanity to ignore one of the conditions.
@@ II.1 Bridge Lemma
To evaluate the conditional distribution for the stochastic variable $x$ under two-time conditions, $\mathbf{P}(x(t)\,|\,x(t_{0}),x(t_{f}))$, we need to relate it to one-time conditioned distributions like $\mathbf{P}(x(t)\,|\,x(t_{0}))$, which we assume given. (And which in our case is provided by Boltzmann’s H-theorem, and its extensions.) We can do this so long as we restrict ourselves to time-symmetric, time-translation invariant one-time conditioned distributions.
@@
Figure 1: An example of a 2LV Boltzmann Bridge in model 1. Many trajectories were sampled with the specified initial and final conditions, for which a sample average was taken. Purple dashed: $\mathbf{E}[S(t)\,|\,S(t_{0}),S(t_{f})].$ Blue dashed: $S_{max}.$ Orange dashed: $S(t_{f})$.
@@ II.2 Marginal Distribution
Returning to the process $x(t)$ defined in Eq. (4), we can ask: what is the marginal $\mathbf{P}(x(t)=a)?$ Using Bayes’ theorem and Eq. (4),
@@
The left hand side of Eq. (8) is independent of $t$ (under the simple assumption that the one-time conditioned distribution is aperiodic and irreducible), thus so is the right hand side, which means the marginal $\mathbf{P}(x(t)=a)$ is time-independent, called stationary.
@@ III.1 Model 1
where $M_{ij}:=\mathbf{P}(r(t+1)=i\,|\,r(t)=j)$ is the $L\times L$ transition matrix, which is time-independent and vanishes if $\,|\,i-j\,|\,>1$. Iterating Eq. (12) allows one to compute the needed one-time conditioned distributions $\mathbf{P}(r(t+k)\,|\,r(t))$, as seen in Fig. 2.
@@
This is the same result as one gets from iterating Eq. (12) in the infinite limit. To compute the two-time condition expectation value of the entropy, we take the $\log$ moment of Eq. (7):
@@
In model 1, the largest macrostate is not necessarily exponentially larger than the other large states. Like any other finite system, the expected value of the entropy of the stationary distribution $\mathbb{E}[S(t)]$ (time-independent), which is what the entropy grows to as $t\rightarrow\infty$, isn’t the exactly entropy of the largest state $r=L$, $S_{max}$. The two entropies agree as $L$ goes to infinity, however there are interesting features of both two and one-time conditioned entropies that depend on the existence of this entropy “gap” between $\mathbb{E}[S(t)]$ and $S_{max}$ as seen in Fig. 6. In the case of the one-time conditioned entropy, the relaxation curve $K(t,a)$ is decreasing if $\log w(a)$ is in the entropy gap - except for an intriguing bump near $t_{0}$. Surprisingly, $K$ does not need to be monotonic in general, even though $R$ is.
Fig. 3 show examples of $\mathbf{E}[S(t)\,|\,r(0)=0,r(t_{f})=b]$ for various values of $b$. What is striking is that, upon close inspection, the concavity $\frac{d^{2}}{dt^{2}}\mathbf{E}[S(t)\,|\,r(0),r(t_{f})]$ can change. There are even extreme cases in which the concavity and/or derivative can change sign twice, as seen in Fig. 4.
@@
We argued in Sec. I that the 2LV phenomenon can intuitively be seen to be true for systems in which $t_{f}-t_{0}$ is much greater than the relaxation time of the system, given that the final condition is below its relaxation value. Two-time conditioned entropies that are conditioned on a final value inside the entropy gap will be concave up near $t_{f}$ if $t_{f}-t_{0}$ is long enough, in complete analogy to how 2LV curves function. Fig. 6 shows the two-time conditioned entropy “tracking” the relaxation curve until near $t_{f}$, when it departs to reach the required final condition (in this scenario and the corresponding 2LV scenario, as $t_{f}\rightarrow\infty$, the time it takes from the entropy to depart from the relaxation time to achieve the final condition becomes only a function of the final condition.) The concavity changes of the two-time conditioned entropy in Fig. 3 however are for $\log w(r(t_{f}))<K(t_{f},0)$, and so cannot be understood in terms of the entropy gap, discussed above, but the concavity change in Fig. 6 is exactly due to the entropy gap. Both of these phenomena will show up in model 2.
Among other strange behaviors that model 1 has, $K(t,a)$ the two-time conditioned entropy can cross the relaxation curve, as seen in Fig. 5. Understanding some of these behaviors is non-obvious, and we will return to them in Section IV.
@@ III.2 Model 2
Consider a box of $N$ particles partitioned into left and right, with a small opening in between the two sides that allows particles to move. The system evolves in discrete time, and at each timestep, each particle has a probability $\alpha_{1}$ of switching sides and probability $\beta_{1}$ of staying on its side. The microstates are each possible configuration (left or right, for each particle) of distinguishable particles, and the macrostates are defined in terms of the macroscopic variable $n_{L}$, the number of particles in the left chamber. (We could take $n_{R}$; it is convention.) As a highly simplified model, we don’t consider motion or volume and thus do not compute the thermodynamic entropy, which depends on such quantities. As in model 1, the Boltzmann entropy suffices for our study:
@@
The marginal distribution, which is necessarily the stationary distribution, can be computed via $\lim_{k\to\infty}\alpha_{k},\beta_{k}=1/2$ to be
@@
The relaxation curve grows to its expected value under the stationary distribution, Eq. (17), which is the equilibrium entropy Eq. (15) only in the limit $N\rightarrow\infty$. This is the same effect as discussed in Sec. III.1. However, the multiplicity $\binom{N}{N/2}$ of microstates for the largest macrostate grows much more quickly as a function of $N$ than the corresponding quantity for model 1, $w(L)$, does, which means that the entropy gap closes more rapidly in model 2 than in model 1 as the systems get larger.
@@ IV Concavity Changes
In this section, we will prove that given some simple assumptions, the two-time conditioned entropy has a change in concavity. We start with two definitions: define $h:=(S(t_{f})-S(t_{0}))/(t_{f}-t_{0})$, with $S(t_{f})>S(t_{0})$ and define the “Bias” of the process $S(t)$ given an initial and final condition to be
@@
As shown in Fig. 4, concavity changes can occur without the conditions $B(t_{0})>h$ and $B(t_{f})>h$ holding, so these conditions should be seen as a sufficient but not necessary condition for the concavity changes.
@@ V Does $\delta=0$ provide a correction?
Define $\delta$ as the offset of the final condition from the relaxation curve for $X$: $\delta:=X(t_{f})-R(t,X(t_{0}))$. It’s then natural to ask if making the final condition live exactly on its relaxation curve affects two time conditional distribution for the entropy:
@@
While not ruling out an analytic solution, we can answer via numerics that they are not equal, as seen in Fig. 9, but the effect disappears in the limit $t_{f}\rightarrow\infty$ as the relaxation curve approaches its asymptote.
@@ VI Open Questions and Future Work
We will conclude with a discussion of open questions and future lines of thinking. The open questions probe specific violations and attenuations of expected 2nd law behavior through the geometry of the two-time condiitoned entropies and through induced probability distributions on $\delta$.
@@
Beyond the open questions, we are interested in studying Boltzmann Bridges in cosmology. In particular, although the thermodynamics of matter is non-equilibrium, the maximum entropy of matter in the universe is increasing [9]. Much more in-depth analysis needs to be done to have more realistic modeling of the universe - see [10] for a study of theoretical predictions of the universe’s entropy over time. Though it might be possible to do such analysis, almost surely in the limit of a large number of particles and a large number of degrees of freedom, the set of values of $\delta$ under which the second law holds has probability infinitesimally close to 1. Despite this, we have raised a careful correction to the standard formulations of the 2nd law from using past hypothesis. This correction, though, can be quite large for the smaller systems we study in this paper, which might have relevance to real microscopic and mesoscopic systems.

### v2

@@ Abstract
It is often stated that the second law of thermodynamics follows from the condition that at some given time in the past the entropy was lower than it is now. Formally, this condition is the statement that $E[S(t)|S(t_{0})]$, the expected entropy of the universe at the current time $t$ conditioned on its value $S(t_{0})$ at a time $t_{0}$ in the past, is an increasing function of $t$. We point out that in general this is incorrect. The epistemic axioms underlying probability theory say that we should condition expectations on all that we know, and on nothing that we do not know. Arguably, we know the value of the universe’s entropy at the present time $t$ at least as well as its value at a time in the past, $t_{0}$. However, as we show here, conditioning expected entropy on its value at two times rather than one radically changes its dynamics, resulting in a unexpected, very rich structure. For example, the expectation value conditioned on two times can have a maximum at an intermediate time between $t_{0}$ and $t$, i.e., in our past. Moreover, it can have a negative rather than positive time derivative at the present. In such “Boltzmann bridge” situations, the second law would not hold at the present time. We illustrate and investigate these phenomena for a random walk model and an idealized gas model, and conclude with open questions about the models.
@@ I Introduction
The seeming paradox of how the second law of thermodynamics can hold despite the time-symmetry of the laws of physics has been a major subject of debate for well over a century [1, 2, 3, 4, 5]. Often the paradox is formalized as the question of how the second law of thermodynamics can apparently hold despite (or due to) the time-symmetry of Boltzmann’s H-theorem. A common claim is that when formulated this way, the paradox is fully resolved by the fact that past entropy was lower in the past than it is now (the “past hypothesis”) [6, 7, 8, 9].
However, this proposed resolution of the paradox is rarely, if ever, presented in a fully formal manner. In this paper we correct this oversight. Since the H-theorem is ultimately a theorem concerning a stochastic process, here we use the theory of Markov processes to investigate it in a fully formal manner. We prove that the proposed resolution of the paradox in terms of the past hypothesis is wrong as stated, and so is misleading without careful qualifications.
Going further, our investigation reveals unforeseen, very rich behavior in the stochastic dynamics of (Boltzmann) entropy through time. Specifically, consider a statistical system, with entropy $S(t)$ evolving in time. Boltzmann’s H-theorem can be summarized as the statement that the conditional expected value of future entropy given current entropy is an increasing function, i.e., for all $k>0$,
@@
$\frac{d}{dt}\mathbb{E}\left[S(t)\,|\,S(t_{0})\right]>0.$ (3)
@@
Accordingly, if we are given the fact that in our universe, $S(t_{0})$ for a time in our past was sufficiently smaller than its maximal value, then the time-derivative of the expected entropy of the universe for all later times, conditioned on that previous value of entropy and nothing else, is non-negative. This is a fully formal statement of the common justification that the second law follows from the past hypothesis.
@@
In this paper we investigate the $t$-dependence of the two-time conditioned expected value of the entropy. In Sec. II.1 we show how to express the two-time conditioned expected entropy in terms of two different one-time conditioned expected entropies. This gives a particularly simple formula for two-time conditioned expected entropy whenever the one-time conditioned expected entropy is time-symmetric — as it is in the time-symmetric version of Boltzmann’s H-theorem in Eq. (2).
We then apply this result to analyze the two-time conditioned expected entropy in toy models of the dynamics of the entropy of the universe. The first of these is a random walk model for the evolution of the state of the universe, and the second is a gas randomly diffusing between two boxes.
We define “2nd law violating” curves (2LV) as solutions to $S(t)$ which have a maximum before $t_{f}$. Intuitively, the 2LV phenomenon is true if $t$ is longer than the relaxation time of the system, or if $S_{now}-S_{past}$ is less that the expected value of the entropy increase (given $S_{past}$ alone) in the $t$ interval.
These observations suggest that the interpretation of the second law as a simple consequence of a past hypothesis requires, at the very minimum, additional careful qualifications. If we use probability theory properly, conditioning on what we know and nothing else, then the past hypothesis that $S(t_{0})$ was low does not establish that the second law holds in the present. Indeed, when probability theory is used correctly, a wealth of rich mathematics is revealed.
In Sec. III.1 and Sec. III.2, we show that both models have the potential to exhibit changes in the concavity and derivative of $\mathbb{E}\left[S(t)\,|\,S(t_{0}),S(t_{f})\right]$, in stark contrast to the familiar behavior. In the first model, this change in concavity common among the space of model parameters. However, in the second model, there is a stronger regularity between 2LV and non-2LV curves that might let us compute the approximate conditions on model parameters that delineate between 2LV and non-2LV behavior. In both models, the two-time conditioned entropy can cross the relaxation curve, which is surprising.
In Sec. IV and Sec. V we investigate the concavity changes in more detail in addition to showing that setting the final condition to be what is expected from the one-time conditioned distribution alone does affect the two-time conditioned distribution. Lastly, in Sec. VI, we outline open questions about the above models and discuss the relevance of our observations for our understanding of the nature of the second law in our universe.
@@ II From physics to distributions
We start from a discrete-valued, discrete time Markov process $x$ taking values on a real set $I$ that has a time-symmetric and time-translation invariant one-time conditioned distribution:
@@
Generally, we want to compute the conditional probability of the stochastic process $x$ conditioned on 2 times: $\mathbf{P}(x(t)\,|\,x(t_{0}),x(t_{f})),$ with $t_{0}<t<t_{f}$. Critically, we seek time-dependence of expectation values in between the initial and final conditions, otherwise we could simply use Markovanity to ignore one of the conditions.
@@ II.1 Bridge Lemma
To evaluate the conditional distribution for the stochastic variable $x$ under two-time conditions, $\mathbf{P}(x(t)\,|\,x(t_{0}),x(t_{f}))$, we need to relate it to one-time conditioned distributions like $\mathbf{P}(x(t)\,|\,x(t_{0}))$, which we assume given. (Which in our case are provided by Boltzmann’s H-theorem and its extensions.) We can do this so long as we restrict ourselves to time-symmetric, time-translation invariant one-time conditioned distributions.
@@
Figure 1: An example of a 2LV Boltzmann Bridge in model 1. Many trajectories were sampled with the specified initial and final conditions, and a sample average was taken. Purple dashed: $\mathbf{E}[S(t)\,|\,S(t_{0}),S(t_{f})].$ Blue dashed: $S_{max}.$ Orange dashed: $S(t_{f})$.
@@ II.2 Marginal Distribution
Returning to the process $x(t)$ defined in Eq. (4), what is the marginal $\mathbf{P}(x(t)=a)?$ Using Bayes’ theorem and Eq. (4),
@@
The left hand side of Eq. (8) is independent of $t$ (under the simple assumption that the one-time conditioned distribution is aperiodic and irreducible),so the right hand side also is, which means the marginal $\mathbf{P}(x(t)=a)$ is time-independent, called stationary.
@@ III.1 Model 1
where $M_{ij}:=\mathbf{P}(r(t+1)=i\,|\,r(t)=j)$ is the time-independent $L\times L$ transition matrix, which vanishes if $\,|\,i-j\,|\,>1$. Iterating Eq. (12) computes the needed one-time conditioned distributions $\mathbf{P}(r(t+k)\,|\,r(t))$, as seen in Fig. 2.
@@
Iterating Eq. (12) in the infinite limit gets this same result. To compute the two-time condition expectation value of the entropy, we take the $\log$ moment of Eq. (7):
@@
In model 1, the largest macrostate is not necessarily exponentially larger than the other large states. Like any other finite system, the expected value of the entropy of the stationary distribution $\mathbb{E}[S(t)]$ (time-independent), which is what the entropy grows to as $t\rightarrow\infty$, isn’t the exactly entropy of the largest state $r=L$, $S_{max}$. The two entropies agree as $L$ goes to infinity, however there are interesting features of both two and one-time conditioned entropies that depend on this entropy “gap” between $\mathbb{E}[S(t)]$ and $S_{max}$ as seen in Fig. 6. In the case of the one-time conditioned entropy, the relaxation curve $K(t,a)$ is decreasing if $\log w(a)$ is in the entropy gap - except for an intriguing bump near $t_{0}$. Surprisingly, $K$ does not need to be monotonic in general, even though $R$ is.
Fig. 3 show examples of $\mathbf{E}[S(t)\,|\,r(0)=0,r(t_{f})=b]$ for various values of $b$. It is striking that (upon close inspection) the concavity $\frac{d^{2}}{dt^{2}}\mathbf{E}[S(t)\,|\,r(0),r(t_{f})]$ can change. There are even extreme cases in which the concavity and/or derivative can change sign twice, as seen in Fig. 4.
@@
We argued in Sec. I that the 2LV phenomenon can intuitively be seen to be true for systems in which $t_{f}-t_{0}$ is much greater than the relaxation time of the system, given that the final condition is below its relaxation value. Two-time conditioned entropies with a final value inside the entropy gap will be concave up near $t_{f}$ if $t_{f}-t_{0}$ is long enough, in complete analogy to how 2LV curves function. Fig. 6 shows the two-time conditioned entropy “tracking” the relaxation curve until near $t_{f}$, when it departs to reach the required final condition (in this scenario and the corresponding 2LV scenario, as $t_{f}\rightarrow\infty$, the time it takes from the entropy to depart from the relaxation time to achieve the final condition becomes only a function of the final condition.) However, the concavity changes of the two-time conditioned entropy in Fig. 3 are for $\log w(r(t_{f}))<K(t_{f},0)$, so they cannot be understood in terms of the entropy gap, discussed above, but the concavity change in Fig. 6 is exactly due to the entropy gap. Both of these phenomena will show up in model 2.
Among other strange behaviors that model 1 has, the two-time conditioned entropy can cross the relaxation curve $K(t,a)$, as seen in Fig. 5. Understanding some of these behaviors is non-obvious, and we will return to them in Sec. IV.
@@ III.2 Model 2
Consider a box of $N$ particles partitioned into left and right, with a small opening in between the two sides that allows particles to switch sides. The system evolves in discrete time, and at each timestep each particle has a probability $\alpha_{1}$ of switching sides and a probability $\beta_{1}$ of staying. The microstates are each possible configuration (left or right, for each particle) of distinguishable particles, and the macrostates are defined in terms of the macroscopic variable $n_{L}$, the number of particles in the left chamber. (We could take $n_{R}$; it is convention.) As a highly simplified model, we don’t consider motion or volume and thus do not compute the thermodynamic entropy, which depends on such quantities. As in model 1, the Boltzmann entropy suffices for our study:
@@
The marginal (stationary) distribution is computed via $\lim_{k\to\infty}\alpha_{k},\beta_{k}=1/2$ to be
@@
The relaxation curve grows to its expected value under the stationary distribution, Eq. (17), which is the equilibrium entropy Eq. (15) only in the limit $N\rightarrow\infty$. This is the same effect as discussed in Sec. III.1. However, the multiplicity $\binom{N}{N/2}$ of microstates for the largest macrostate grows much more quickly as a function of $N$ than the corresponding quantity for model 1, $w(L)$, does. T his means that the entropy gap closes more rapidly in model 2 than in model 1 as the systems get larger.
@@ IV Concavity Changes
In this section, we will prove that the two-time conditioned entropy has a change in concavity under some simple assumptions,. We start with two definitions: define $h:=(S(t_{f})-S(t_{0}))/(t_{f}-t_{0})$, with $S(t_{f})>S(t_{0})$ and define the “Bias” of the process $S(t)$ given an initial and final condition to be
@@
As shown in Fig. 4, concavity changes can occur without the conditions $B(t_{0})>h$ and $B(t_{f})>h$ holding, so these conditions should be seen as a sufficient but not necessary.
@@ V Does $\delta=0$ provide a correction?
Define $\delta$ as the offset of the final condition from the relaxation curve for $X$: $\delta:=X(t_{f})-R(t,X(t_{0}))$. It’s natural to ask if making the final condition live exactly on its relaxation curve affects two time conditional distribution for the entropy:
@@
While there might be an analytic solution, we can answer via numerics that they are not equal, as seen in Fig. 9, but the effect disappears in the limit $t_{f}\rightarrow\infty$ as the relaxation curve approaches its asymptote.
@@ VI Open Questions and Future Work
Our analysis raises many open question about specific violations and changes of expected 2nd law behavior:
@@
Beyond the open questions, we are interested in studying Boltzmann Bridges in cosmology. In particular, although the thermodynamics of matter is non-equilibrium, the maximum entropy of matter in the universe is increasing [14]. Much more in-depth analysis needs to be done to have more realistic modeling of the universe - see [15] for a study of theoretical predictions of the universe’s entropy over time. Though such analysis might be possible, almost surely in the limit of a large number of particles and a large number of degrees of freedom, the set of values of $\delta$ under which the second law holds has probability infinitesimally close to 1. Despite this, we have raised a careful correction to the standard formulations of the 2nd law from using past hypothesis. The correction can be quite large for the smaller systems we study in this paper, which might have relevance to real microscopic and mesoscopic systems.